# MrT5: Dynamic Token Merging for Efficient Byte-level Language Models

### v2

@@ 1 Introduction
Subword tokenization, typically via algorithms such as byte-pair encoding (Sennrich et al., 2016) or SentencePiece (Kudo & Richardson, 2018), is a fundamental text preprocessing step that has become ubiquitous in modern large language models (LLMs). Subword tokenizers divide text into meaningful units known as tokens, which closely resemble words or parts of words. Tokenization can be seen as a form of compression, since it reduces the sequence length of the input passed to the compute-intensive Transformer (Vaswani et al., 2017). However, subword tokenizers have several drawbacks. For example, they are not very robust to character-level noise and manipulations, such as spelling errors (Kaushal & Mahowald, 2022; Huang et al., 2023); they directly impact how models process digits and perform arithmetic (Singh & Strouse, 2024); and they have disproportionate compression rates for different languages and scripts (Ahia et al., 2023; Petrov et al., 2023). In addition, current language model APIs charge users per-token, and such discrepancies can cause users of certain languages to be overcharged due to poorer compression. [^1] See also Andrej Karpathy’s tweets on tokenization: https://x.com/karpathy/status/1759996551378940395; https://x.com/karpathy/status/1657949234535211009
@@ 2 Background: Tokenization-Free Models
MegaByte (Yu et al., 2023) scales byte-level decoders to long-context tasks by segmenting sequences into fixed-length “patches,” though these may not align with meaningful units of text. SpaceByte (Slagle, 2024) applies global Transformer blocks to specific byte types, such as spaces, but these bytes are also not chosen dynamically. Hourglass Transformers (Nawrot et al., 2022) incorporate hierarchical downsampling and upsampling at different layers within decoder models, and the architecture was later extended with a gradient-based tokenization module that dynamically pools characters using a boundary predictor (Nawrot et al., 2023; Ahia et al., 2024). MANTa (Anagnostidis et al., 2024) also learns token boundaries via sliding window attention. More recently, the Byte Latent Transformer (BLT, Pagnoni et al., 2024) and EvaByte (Zheng et al., 2025) have demonstrated that tokenizer-free models can achieve performance comparable to subword LLMs. BLT utilizes a dynamic boundary predictor to segment sequences into variably sized patches, leveraging byte entropies derived from an external byte-level language model to determine patch boundaries. EvaByte uses multibyte prediction and an efficient linear attention mechanism to improve scalability and decoding speed.
@@
This paper focuses on ByT5 (Xue et al., 2022), a byte-level sequence-to-sequence Transformer architecture that serves as a counterpart to mT5 (Xue et al., 2021), the multilingual version of T5 (Raffel et al., 2020). Unlike the models previously discussed, ByT5 uses no soft tokenization or downsampling steps. To compensate for the loss of the parameters that would normally be used for subword embeddings, ByT5 has a “heavy” encoder with a larger number of layers than the decoder. While ByT5 shows impressive performance on a variety of downstream tasks, its heavy encoder, large model and feed forward dimensionalities, and short input sequence length (1024 bytes) make it quite inefficient. The architecture itself requires about 1.2 times more operations than mT5, which contributes to a 33% longer pre-training wall clock time for ByT5, even though it was trained on only a quarter of the data used for mT5. In terms of inference speed on downstream tasks, ByT5 can be up to 10 times slower than mT5, depending on the input sequence length.
MrT5 modifies the ByT5 architecture to make it more efficient. Unlike previous work, MrT5’s deletion gating mechanism can be added to a pre-trained model like ByT5 with minimal fine-tuning using a small number of extra parameters. The gating can also be applied to models trained from scratch. For further related work on adjacent topics like early exit models, long-context, and additional tokenization-free architectures, see Appendix A.
@@ 3 The MrT5 Model Architecture
The unique aspect of MrT5 is its deletion gating mechanism: after a fixed encoder layer, a delete gate determines which tokens in the sequence should be kept to be processed by later layers, and which tokens may be deleted, thereby “merging” information into a shorter sequence. Our choice to delete tokens only in a single layer has three motivations: (1) we want to avoid the overhead of executing the deletion algorithm multiple times; (2) we observe that performance stabilizes after a few initial layers (see Section 7); and (3) in terms of minimizing computation costs, the deletion is most beneficial if done in an early layer.
@@ 3.1 Deletion Gating Mechanism
$\mathbf{G}$ $=k\sigma(\mathbf{H}_{l}\mathbf{W}+\mathbf{1}_{N}b)$ (1)
where $\mathbf{H}_{l}\in\mathbb{R}^{N\times d}$ are the hidden states output by layer $l$; $\mathbf{W}\in\mathbb{R}^{d\times 1}$; $b\in\mathbb{R}$; $\mathbf{G}\in\mathbb{R}^{N\times 1}$; $k$ is a large negative constant; $N$ is the encoder input sequence length; $d$ is ByT5’s hidden state/model dimensionality; and $\mathbf{1}_{N}\in\mathbb{R}^{N\times 1}$ is a vector of ones. The gating activation function is a rescaled and translated sigmoid function, bounded between $k$ and 0. In our experiments, we use $k=-30$. During training, our experiments include adding Gumbel noise to the gate’s logits to encourage exploration of gate values. However, we found that the model performs well even without it, making this step optional. See Appendix B for the Gumbel noise formulas.
@@
During training, MrT5 deletes tokens softly, where the outputs of the gating mechanism $\mathbf{G}$ are applied as a soft attention mask. The outputs are added directly to the self-attention mechanism of the subsequent encoder layers, as well as the cross attentions between the decoder and encoder. As an example, the hidden states of encoder layer $l+1$ are defined as:
$\mathbf{H}_{l+1}$ $=\mathrm{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}+\mathbf{1}_{N}\mathbf{G}^{\top}\right)\mathbf{V}$ (2)
where $\mathbf{Q},\mathbf{K},\mathbf{V},\mathbf{H}_{l+1}\in\mathbb{R}^{N\times d}$ and $\mathbf{1}_{N}\in\mathbb{R}^{N\times 1}$ is a vector of ones. A token at sequence position $i\in[1,N]$ and $\mathbf{G}_{i}\approx 0$ will not be masked, whereas a token at sequence position $j\neq i\in[1,N]$ and $\mathbf{G}_{j}\approx k$ will be masked, since $k$ is a large negative constant. Though soft deletion does not reduce the sequence length, we apply it during training to emulate the effect of token deletion while being fully differentiable. To see efficiency gains during inference, we apply hard deletion, where the hidden states are removed from the sequence, determined by a hard threshold; we set this threshold to be $\frac{k}{2}$, half of the range of the delete gate’s output.
@@ 3.2 Gate Regularizer
Setting $\alpha$ by hand allows the model to dynamically discover the deletion ratio depending on the difficulty of the task. However, we can also optimize for a specific ratio of deleted tokens using an algorithm that resembles the proportional-integral controller (PI-controller) of classical control theory. We additionally use an exponential moving average on the P-term. Let’s call the target deletion ratio $\delta\in[0,1]$, the proportion of deleted tokens in the current batch $\hat{\delta}_{t}\in[0,1]$, and the regularization hyperparameter for the current batch $\alpha_{t}$. We update $\alpha$ as follows:
@@
where $\mathrm{clamp}(x)=\max(x,0)$. We initialize $p_{0}=0$ and $i_{0}=0$. In all of our experiments, we use $\gamma=0.9$. For most experiments (unless otherwise noted), we found that using $k_{p}=0.5$ and $k_{i}=1\mathrm{e}{-5}$ worked well in practice. This method is easier to use than manually setting $\alpha$ and allows $\alpha$ to change dynamically as the model undergoes phase transitions during training, resulting in more stable learning. Besides our diagnostic task models, all MrT5 models are trained with this PI-controller.
@@ 5 Continued Pre-training
We train several MrT5 models with different PI-controller target deletion ratios: $\delta\in\{0.3,0.4,0.5,0.6,0.7\}$. This is a continued pre-training setup; we load the pre-trained weights and use the architecture settings of ByT5 Small (300M parameters) and only randomly initialize the weights of MrT5’s gating mechanism. Based on a sweep of different layers, we place the gating mechanism after encoder layer $l=3$ (see Section 7). In addition to the MrT5 models, we also train several baselines:
@@ 7 Analysis
Figure 4: BPB and inference runtime for a single sequence for MrT5 models with delete gates at different layers ($l\in[1,5]$). All MrT5 models are trained with a PI-controller with a target deletion ratio of $\delta=0.5$. BPB is consistently higher in early layers; since the gate should be placed as early as possible, we select layer 3 as optimal.
@@ Acknowledgments
Steps for reproducing each of our experiments are detailed in Appendix E. Descriptions of the model architectures and training configurations/hyperparameters for our diagnostic task experiments are provided in Section E.1; details of the model architectures, span corruption data preprocessing steps, and training configurations/hyperparameters for continued pre-training are provided in Section E.2; and training configurations/hyperparameters for fine-tuning on the XNLI and character-level downstream tasks are provided in Section E.3. We provide our source code at https://github.com/jkallini/mrt5.
@@ Appendix A Further Related Work
Other work has attempted to address issues with long sequences in Transformer models more generally. For example, Perceiver AR (Hawthorne et al., 2022) is an autoregressive model that employs cross-attention to condense inputs into a smaller set of latent representations. This approach enables it to handle attention over large input sequences of up to a hundred thousand tokens. However, the authors did not specifically evaluate the model on character-level tasks or language tasks without tokenization. Other solutions include Nugget (Qin & Van Durme, 2023; Qin et al., 2023), which encodes the whole sentence, but passes only a dynamic subset of the embeddings to the decoder. This approach does not save compute on the encoder side.
More recently, GemFilter (Shi et al., 2024) has improved inference runtimes in decoder-only models by using early LLM layers as a filter to select input tokens to be processed by the full model; while this method requires no additional training, it requires two forward passes during inference.
@@
Before the Transformer, character-level LSTMs, with their compact vocabularies, performed well on a range of NLP tasks and were particularly effective for languages with complex morphology (Gillick et al., 2016; Kim et al., 2016; Ballesteros et al., 2015; Cherry et al., 2018). More recently, MambaByte (Wang et al., 2024) has provided a tokenization-free alternative to the Mamba architecture (Gu & Dao, 2023), trained directly on byte sequences. Tokenization-free methods have also been applied to model other modalities, such as speech (Irie et al., 2017) or text rendered as images (Rust et al., 2023).
@@ Appendix C Theoretical Compute Savings
Note that $\delta=0$ corresponds to the baseline ByT5. For MrT5 Small, $d_{\text{model}}=1472,d_{\text{ff}}=3584,L_{E}=12,L_{D}=4,l_{\mathbf{G}}=3,N_{E}=1024,N_{D}=189$. Given these numbers, we plot the total reduction in compute (compared to ByT5) as a function of $\delta$ in Fig. 5.
@@ E.2 Continued Pre-training Details
All MrT5 and baseline models use the model configuration of a standard ByT5 Small, which has $d_{\text{ff}}=3584$, $d_{\text{model}}=1472$, 12 encoder layers, 4 decoder layers, 6 attention heads in each layer, and 300M total parameters. MrT5’s gating mechanism introduces an additional 3000 parameters; the boundary predictor baseline’s downsampling module introduces an additional 5M parameters; and the convolutional pooling baseline’s downsampling module introduces and additional 6.5M parameters. All models use $\mathrm{softmax}_{1}$ in their attention mechanisms.
@@ E.3 Downstream Task Details
For all downstream tasks, we fine-tune all MrT5 and baseline models using the AdamW optimizer with a cosine learning rate schedule and no warmup. MrT5 models use a PI-controller to achieve a target deletion ratio and use attention score regularization, as described in Appendix D. Training and evaluation details for each task are summarized in Table 6.
@@
Task Steps Epochs Batch Size Max Seq. Length LR PI-Controller
@@ F.2 Description of Boundary Predictor Baselines
Once boundaries $b_{t}$ are determined, tokens within the same segment are aggregated using mean pooling into a shortened sequence length $S=1+\sum_{t}b_{t}$. This is computed via a binary assignment matrix $\mathbf{B}\in\mathbb{R}^{N\times S}$, where $\mathbf{B}_{ij}=1$ if token $i$ is assigned to segment $j$, and 00 otherwise. The pooled segment representations $\mathbf{S}\in\mathbb{R}^{S\times d}$ (where $d$ is the model dimensionality) are computed via weighted averaging of the hidden states $\mathbf{H}\in\mathbb{R}^{N\times d}$:
$\mathbf{S}=(\mathbf{B}^{\top}\mathbf{H})\oslash(\mathbf{B}^{\top}\mathbf{1}_{N\times d})$ (19)

### v3

@@ 1 Introduction
Subword tokenization, typically via algorithms such as byte-pair encoding (Sennrich et al., 2016) or SentencePiece (Kudo & Richardson, 2018), is a fundamental text preprocessing step that has become ubiquitous in modern large language models (LLMs). Subword tokenizers divide text into meaningful units known as tokens, which closely resemble words or parts of words. Tokenization can be seen as a form of compression, since it reduces the sequence length of the input passed to the compute-intensive Transformer (Vaswani et al., 2017). However, subword tokenizers have several drawbacks. For example, they are not very robust to character-level noise and manipulations, such as spelling errors (Kaushal & Mahowald, 2022; Huang et al., 2023); they directly impact how models process digits and perform arithmetic (Singh & Strouse, 2024; Zhou et al., 2024); and they have disproportionate compression rates for different languages and scripts (Ahia et al., 2023; Petrov et al., 2023). In addition, current language model APIs charge users per-token, and such discrepancies can cause users of certain languages to be overcharged due to poorer compression. [^1] See also Andrej Karpathy’s tweets on tokenization: https://x.com/karpathy/status/1759996551378940395; https://x.com/karpathy/status/1657949234535211009
@@ 2 Background: Tokenization-Free Models
MegaByte (Yu et al., 2023) scales byte-level decoders to long-context tasks by segmenting sequences into fixed-length “patches,” though these may not align with meaningful units of text. SpaceByte (Slagle, 2024) applies global Transformer blocks to specific byte types, such as spaces, but these bytes are also not chosen dynamically. Hourglass Transformers (Nawrot et al., 2022) incorporate hierarchical downsampling and upsampling at different layers within decoder models, and the architecture was later extended with a gradient-based tokenization module that dynamically pools characters using a boundary predictor (Nawrot et al., 2023; Ahia et al., 2024). MANTa (Anagnostidis et al., 2024) also learns token boundaries via sliding window attention. More recently, the Byte Latent Transformer (BLT, Pagnoni et al., 2024) and EvaByte (Zheng et al., 2025) have shown that tokenizer-free models can scale more efficiently than subword LLMs. BLT utilizes a dynamic boundary predictor based on byte entropies to segment sequences into variably sized patches, and EvaByte uses multibyte prediction and linear attention to improve scalability and decoding speed.
@@
This paper focuses on ByT5 (Xue et al., 2022), a byte-level sequence-to-sequence Transformer architecture designed as a counterpart to mT5 (Xue et al., 2021), the multilingual extension of T5 (Raffel et al., 2020). Unlike the models discussed previously, ByT5 uses no soft tokenization or downsampling steps to reduce the sequence length. To compensate for the loss of the parameters that would normally be used for subword embeddings, ByT5 has a “heavy” encoder with more layers than the decoder. While ByT5 matches or outperforms mT5 on a variety of downstream tasks, its heavy encoder, large model and feed forward dimensionalities, and short input sequence length (1024 bytes) make it quite inefficient. ByT5 requires 1.2 times more operations than mT5, and it can be up to 10 times slower at inference, depending on the encoder’s input length.
MrT5 modifies the ByT5 architecture to make it more efficient. Unlike previous work, MrT5’s deletion gating mechanism can be applied to a pre-trained model with minimal fine-tuning and few additional parameters, or to models trained from scratch. For further related work on adjacent topics like early exit models, long-context, and additional tokenization-free architectures, see Appendix A.
@@ 3 The MrT5 Model Architecture
MrT5 introduces a unique deletion gating mechanism: in a fixed encoder layer, a delete gate selects which tokens to keep for further processing and which to discard, effectively merging information into a shorter sequence. We apply this gating at a single layer for three reasons: (1) to avoid the overhead of executing the deletion algorithm multiple times, (2) because performance stabilizes after early layers (Section 7), and (3) because early deletion yields the greatest computational savings.
@@ 3.1 Deletion Gating Mechanism
$\mathbf{G}$ $=k\sigma(\mathrm{LayerNorm}(\mathbf{H}_{l})\mathbf{W}+\mathbf{1}_{N}b)$ (1)
where $\mathbf{H}_{l}\in\mathbb{R}^{N\times d_{\text{model}}}$ are the hidden states output by layer $l$; $\mathbf{W}\in\mathbb{R}^{d_{\text{model}}\times 1}$; $b\in\mathbb{R}$; $\mathbf{G}\in\mathbb{R}^{N\times 1}$; $k$ is a large negative constant; $N$ is the encoder input sequence length; $d_{\text{model}}$ is ByT5’s hidden state/model dimensionality; and $\mathbf{1}_{N}\in\mathbb{R}^{N\times 1}$ is a vector of ones. $\mathrm{LayerNorm}(\mathbf{H}_{l})$ denotes the application of layer normalization. Following the T5 architecture, this is implemented as root mean square layer normalization (RMSNorm, Zhang & Sennrich, 2019). The gating activation function is a rescaled and translated sigmoid function, bounded between $k$ and 0. In our experiments, we use $k=-30$. During training, our experiments include adding Gumbel noise to the gate’s logits to encourage exploration of gate values. However, we found that the model performs well even without it, making this step optional. See Appendix B for the Gumbel noise formulas.
MrT5’s delete gate introduces only $12d_{\text{model}}+$ 1 additional parameters in total; $d_{\text{model}}$ for the weight vector $\mathbf{W}$, $d_{\text{model}}$ for the layer normalization, and one for the bias term. This makes the method highly parameter-efficient.
@@
During training, MrT5 deletes tokens softly, where the outputs of the gating mechanism $\mathbf{G}$ are applied as a soft attention mask. These outputs are added directly to the self-attention mechanism of the subsequent encoder layers, as well as to the cross-attention layers between the decoder and encoder. The modified attention mechanism for each attention head is defined as:
$\mathrm{SoftDeletionAttention}(\mathbf{Q},\mathbf{K},\mathbf{V})$ $=\mathrm{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{\text{head}}}}+\mathbf{1}_{N}\mathbf{G}^{\top}\right)\mathbf{V}$ (2)
where $d_{\text{head}}$ is the dimensionality of the query, key, and value vectors for a single attention head; $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times d_{\text{head}}}$; and $\mathbf{1}_{N}\in\mathbb{R}^{N\times 1}$ is a vector of ones. A token at sequence position $i\in[1,N]$ with $\mathbf{G}_{i}\approx 0$ will not be masked, whereas a token at position $j\neq i\in[1,N]$ with $\mathbf{G}_{j}\approx k$ will be masked, since $k$ is a large negative constant. Though soft deletion does not reduce the sequence length, we apply it during training to emulate the effect of token deletion while being fully differentiable. To see efficiency gains during inference, we apply hard deletion, where the hidden states are removed from the sequence, determined by a hard threshold; we set this threshold to be $\frac{k}{2}$, half of the range of the delete gate’s output.
@@ 3.2 Gate Regularizer
Setting $\alpha$ by hand allows the model to dynamically discover the deletion ratio depending on the difficulty of the task. However, we can also optimize for a specific ratio of deleted tokens using an algorithm that resembles the proportional-integral controller (PI controller) of classical control theory. We additionally use an exponential moving average on the P-term. Let’s call the target deletion ratio $\delta\in[0,1]$, the proportion of deleted tokens in the current batch $\hat{\delta}_{t}\in[0,1]$, and the regularization hyperparameter for the current batch $\alpha_{t}$. We update $\alpha$ as follows:
@@
where $\mathrm{clamp}(x)=\max(x,0)$. We initialize $p_{0}=0$ and $i_{0}=0$. In all of our experiments, we use $\gamma=0.9$. For most experiments (unless otherwise noted), we found that using $k_{p}=0.5$ and $k_{i}=1\mathrm{e}{-5}$ worked well in practice. This method is easier to use than manually setting $\alpha$ and allows $\alpha$ to change dynamically as the model undergoes phase transitions during training, resulting in more stable learning. Besides our diagnostic task models, all MrT5 models are trained with this PI controller.
@@ 5 Continued Pre-training
We train several MrT5 models with different PI controller target deletion ratios: $\delta\in\{0.3,0.4,0.5,0.6,0.7\}$. This is a continued pre-training setup; we load the pre-trained weights and use the architecture settings of ByT5 Small (300M parameters) and only randomly initialize the weights of MrT5’s gating mechanism. Based on a sweep of different layers, we place the gating mechanism after encoder layer $l=3$ (see Section 7). In addition to the MrT5 models, we also train several baselines:
@@ 7 Analysis
Figure 4: BPB and inference runtime for a single sequence for MrT5 models with delete gates at different layers ($l\in[1,5]$). All MrT5 models are trained with a PI controller with a target deletion ratio of $\delta=0.5$. BPB is consistently higher in early layers; since the gate should be placed as early as possible, we select layer 3 as optimal.
@@ Acknowledgments
Steps for reproducing each of our experiments are detailed in Appendix E. Descriptions of the model architectures and training configurations/hyperparameters for our diagnostic task experiments are provided in Section E.1; details of the model architectures, span corruption data preprocessing steps, and training configurations/hyperparameters for continued pre-training are provided in Section E.2; and training configurations/hyperparameters for fine-tuning on the multilingual and character-level downstream tasks are provided in Section E.3. We provide our source code at https://github.com/jkallini/mrt5.
@@ Appendix A Further Related Work
Other work has addressed long-context modeling more broadly. Perceiver AR (Hawthorne et al., 2022) uses cross-attention to map inputs to a smaller set of latent vectors, enabling autoregressive modeling over sequences up to 100K tokens. However, it was not evaluated on character-level tasks or language tasks without tokenization. Other solutions include Nugget (Qin & Van Durme, 2023; Qin et al., 2023), which encodes the whole sentence, but passes only a dynamic subset of the embeddings to the decoder. This approach does not save compute on the encoder side. Recently, GemFilter (Shi et al., 2024) has improved inference runtimes in decoder-only models by using early LLM layers as a filter to select input tokens to be processed by the full model; while this method requires no additional training, it requires two forward passes during inference.
@@
Before the Transformer, character-level LSTMs, with their compact vocabularies, were effective across a range of NLP tasks, particularly in morphologically rich languages (Gillick et al., 2016; Kim et al., 2016; Ballesteros et al., 2015; Cherry et al., 2018). Building on this idea, CharBERT (Ma et al., 2020) enhances subword models with character-level embeddings from a bidirectional GRU and a noise-aware pretraining task to improve robustness. More recently, MambaByte (Wang et al., 2024) extends the Mamba architecture (Gu & Dao, 2023) to byte-level inputs, offering a fully tokenization-free alternative. Tokenization-free approaches have also been explored in other modalities, including speech (Irie et al., 2017) and text rendered as images (Rust et al., 2023).
@@ Appendix C Theoretical Compute Savings
Note that $\delta=0$ corresponds to the baseline ByT5. For MrT5 Small, $d_{\text{model}}=1472$, $d_{\text{ff}}=3584$, $L_{E}=12$, $L_{D}=4$, $l_{\mathbf{G}}=3$, $N_{E}=1024$, and $N_{D}=189$. Given these numbers, we plot the total reduction in compute (compared to ByT5) as a function of $\delta$ in Fig. 5.
@@ E.2 Continued Pre-training Details
All MrT5 and baseline models use the model configuration of a standard ByT5 Small, which has $d_{\text{ff}}=3584$, $d_{\text{model}}=1472$, 12 encoder layers, 4 decoder layers, 6 attention heads in each layer, and 300M total parameters. MrT5’s gating mechanism introduces an additional 2,945 parameters; the boundary predictor baseline’s downsampling module introduces an additional 5M parameters; and the convolutional pooling baseline’s downsampling module introduces and additional 6.5M parameters. All models use $\mathrm{softmax}_{1}$ in their attention mechanisms.
@@ E.3 Downstream Task Details
For all downstream tasks, we fine-tune all MrT5 and baseline models using the AdamW optimizer with a cosine learning rate schedule and no warmup. MrT5 models use a PI controller to achieve a target deletion ratio and use attention score regularization, as described in Appendix D. Training and evaluation details for each task are summarized in Table 6.
@@
Task Steps Epochs Batch Size Max Seq. Length LR PI Controller
@@ F.2 Description of Boundary Predictor Baselines
Once boundaries $b_{t}$ are determined, tokens within the same segment are aggregated using mean pooling into a shortened sequence length $S=1+\sum_{t}b_{t}$. This is computed via a binary assignment matrix $\mathbf{B}\in\mathbb{R}^{N\times S}$, where $\mathbf{B}_{ij}=1$ if token $i$ is assigned to segment $j$, and 00 otherwise. The pooled segment representations $\mathbf{S}\in\mathbb{R}^{S\times d_{\text{model}}}$ (where $d_{\text{model}}$ is the model dimensionality) are computed via weighted averaging of the hidden states $\mathbf{H}\in\mathbb{R}^{N\times d_{\text{model}}}$:
$\mathbf{S}=(\mathbf{B}^{\top}\mathbf{H})\oslash(\mathbf{B}^{\top}\mathbf{1}_{N\times d_{\text{model}}})$ (19)